39 Under the direct write-off period and is collected in the the journal entry? method, when an acount receivable is written off in one accounting ollowing accounting period, which of the following would be included in an accou per 40. A note on which no rate of interest is specified is a(n) 41. When a business endorses 42. The person who promises 43. The face amount of a note that is promised to be paid at maturity is called the 44, A $7,300, 11.9% note is dated April 21 and is due in 60 days. The amount of interest on the due date a note and transfers it to a bank, the process is called to pay a certain amount of money at a definite future time is called the of the note would be 45. A $6,700, 85% note is dated April 10 and is due in 75 days. The maturity value of the note would be
Answer #1
39) The entry would be:
Accounts receivable
Bad debts recovered
and
Cash
Accounts receivable
40) A Non-interest bearing note
41) Negotiation
42) Promisor
43) Principal
44) Interest = 7300*11.9%*60/365 = $     142.80
45) Maturity value = 6700+6700*8.5%*75/365 = $ 6,817.02